OpenStudy (anonymous):

An object is propelled vertically upward from the top of a 256-foot building. The quadratic function s(t)=-16t^2+144t+256 models the ball's height above the ground, s(t) in feet, t seconds after it was thrown. How many seconds does it take until the object finally hits the ground? Round to the nearest tenth of a second if necessary. a. 1.5 seconds b. 4.5 seconds c. 10.5 seconds d. 2 seconds

OpenStudy (alfie):

Hitting the ground => final Y position = 0 sooo...

OpenStudy (anonymous):

when the object is on the ground the hight is zero. therefore s(t)=0 therefore 16t^2+144t+256=0 t^2+9t+16=0 (dividing by 16) using quadratic formula (-9+-root(9^2-(4x-1x16))/2=t t=(-9+-(root (145))/2 = either 1.52... or -10.5.... it can't be the negative one therefore answer a

hero (hero):

Answer is C

OpenStudy (alfie):

Yup, it's C. Gabriel you messed up a minus sign in front of the sixteen back there :). But good job on getting the problem straight ;)

hero (hero):

Oh well, too late now

OpenStudy (alfie):

too late for what? o_O

hero (hero):

The OP is gone

OpenStudy (alfie):

What does OP stand for? lol (i phail)

hero (hero):

Original Poster

OpenStudy (alfie):

oh well, :)... ! The hope is he got how to do it, and not just how to copy a quadratic equation solution :D...!

OpenStudy (anonymous):

just didn't divide through by minus 16 was all, i only noticed it half way through answering it because of my bad eyesight :p

OpenStudy (anonymous):

i'm here, had to step away for a minute.. so... it's c?

OpenStudy (anonymous):

yep seems so.